\begin{thebibliography}{00}

%1
\bibitem{BHV} L.~Brink, H.~Hansson and M.A.~Vasiliev,
              "Explicit solution to the N-body Calogero problem",
             Phys.~Lett. {\bf B286} (1992) 109-111.

%2
\bibitem{Cal1} F.~Calogero,
``Solution of a three-body problem in one dimension",
J.~Math.~Phys. {\bf 10} (1969) 2191-2196;

\bibitem{Cal2} F.~Calogero,
``Ground state of a one dimensional N-body problem",
J.~Math.~Phys. {\bf 10} (1969) 2197-2200;

\bibitem{Cal3} F.~Calogero,
``Solution of the One-Dimensional N-Body Problems with Quadratic and/or Inversely Quadratic Pair Potentials",
J.~Math.~Phys. {\bf 12} (1971) 419,
             https://doi.org/10.1063/1.1665604.


%3
\bibitem{Dunkl} C.F.~Dunkl,
"Differential-difference operators associated to reflection groups",
Trans. Amer. Math. Soc., {\bf 311} (1) (1989) 167-183,
                          doi:10.2307/2001022.


%4
\bibitem{sra}
P.~Etingof and V.~Ginqzburg, Symplectic reflection algebras,
Calogero--Moser space, and
deformed Harish--Chandra homomorphism,
Inv. Math. 147 (2002), 243--348.



%7
\bibitem{K2} S.E.~Konstein, ``3-particle Calogero Model: Supertraces and Ideals on the Algebra of Observables",
              Teor.~Mat.~Fiz., {\bf 116} (1998) 122; arXiv:hep-th/9803213.


%9
\bibitem{stek} S.E.~Konstein and R.~Stekolshchik,
               \lq\lq Klein operator and the Number of Traces and Supertraces
              on the Superalgebra of Observables of Rational Calogero Model
              based on the Root System",
              Journal of Nonlinear Mathematical Physics,
              Vol. 20, No. 2 (June 2013), 295-308;

%19.
\bibitem{KTroot} S.E.~Konstein and I.V.~Tyutin,
``Traces on the Algebra of Observables of the Rational Calogero Model Based on the
Root System", Journal of Nonlinear Mathematical Physics, 20:2, (2013)
271-294; arXiv:1211.6600


%18.
\bibitem{KTSRA} S.E.~Konstein and I.V.~Tyutin,
``The number of independent traces and supertraces on symplectic
reflection algebras", Journal of Nonlinear Mathematical Physics, 21:3, (2014) 308-335;
arXiv:1308.3190

%17
\bibitem{2m+1} S.E.~Konstein and I.V.~Tyutin,
``Ideals generated by traces or by supertraces in the symplectic
reflection algebra $H_{1,\nu}(I_2(2m+1))$'',
Journal of Nonlinear Mathematical Physics, 24:3, 405-425 (2017),
DOI:10.1080/14029251.2017.1341702; arXiv:1612.00536



%6
\bibitem{KV} S.E.~Konstein and M.A.~Vasiliev,
``Supertraces on the Algebras of Observables of the Rational Calogero Model with Harmonic Potential",
              J.~Math.~Phys. {\bf 37} (1996) 2872.

%12
\bibitem{OP} M.A.~Olshanetsky and A.M.~Perelomov,
             ``Quantum integrable systems related to lie algebras",
              Phys. Rep., {\bf 94}, Issue 6, (1983) 313-404,
              DOI:  10.1016/0370-1573(83)90018-2.

%14
\bibitem{Poly} A.~Polychronakos,
               ``Exchange operator formalism for integrable systems of particles",
               Phys.~Rev.~Lett. {\bf 69} (1992) 703-705.

%15
\bibitem{V1} M.A.~Vasiliev,
``Quantization on sphere and high-spin superalgebras ",
JETP Letters, {\bf 50} (1989) 377-379;

\bibitem{V2} M.A.~Vasiliev,
``Higher spin algebras and quantization on the sphere and hyperboloid",
              Int. J. Mod. Phys. {\bf A6} No. 07 (1991) 1115-1135,
https://doi.org/10.1142/S0217751X91000605.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{euscript,amsmath, amssymb, amsfonts}
\usepackage{color}

\newcommand{\R}{{\cal R}}
\newcommand{\str}{{\sf sp} }
\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bee}{\begin{eqnarray}}
\newcommand{\eee}{\end{eqnarray}}

\newcommand{\vv}{{\vec v}}

\newcounter{theorem}

\makeatletter \@addtoreset{theorem}{section}

\newcounter{corollary}

\makeatletter \@addtoreset{corollary}{section}

\newcounter{lemma}

\makeatletter \@addtoreset{lemma}{section}

\newcounter{proposition}

\makeatletter \@addtoreset{proposition}{section}

\newcounter{conjecture}

\makeatletter \@addtoreset{conjecture}{section}

\newcounter{remark}

\makeatletter \@addtoreset{remark}{section}

\newcounter{definition}

\makeatletter \@addtoreset{definition}{section}

\makeatletter \@addtoreset{equation}{section}

\newenvironment{proof}[1][Proof]{\noindent\textsf{#1.\ }}
{\hfill {\small $\square$}}

\begin{document}

\sloppy \title
 {
Connection between the ideals generated by traces and by supertraces
in the superalgebras of
observables of Calogero models
}

\author
 {
 S.E. Konstein%
\thanks{ I.E. Tamm Department of Theoretical Physics,
          P.N. Lebedev Physical Institute, RAS
          119991, Leninsky prosp., 53, Moscow, Russia}
\thanks{E-mail: konstein@lpi.ru}
 ,
 I.V. Tyutin$^*$%
\thanks{E-mail: tyutin@lpi.ru}     }

\date{
}

\maketitle
\thispagestyle{empty}

\begin{abstract}

If $G$ is a finite Coxeter group, then symplectic reflection algebra
$H:=H_{1,\eta}(G)$ has Lie algebra $\mathfrak {sl}_2$ of inner derivations and can be decomposed
under spin: $H=H_0 \oplus H_{1/2} \oplus H_{1} \oplus H_{3/2} \oplus ...$.
We show that if the ideals $\mathcal I_i$ ($i=1,2$) of all the vectors
from the kernel
of degenerate bilinear forms $B_i(x,y):=sp_i(x\cdot y)$,
where $sp_i$ are (super)traces on $H$, do exist, then $\mathcal I_1=\mathcal I_2$
if and only if $\mathcal I_1 \bigcap H_0=\mathcal I_2 \bigcap H_0$.

\end{abstract}


\section{Preliminaries and notation}

Let ${\mathcal A}$ be an associative superalgebra with parity $\pi$.
All expressions of linear algebra are given for homogenous elements only
and are supposed to be extended to inhomogeneous elements via linearity.

\begin{definition}\label{str}
A linear function $str$ on ${\mathcal A}$ is called a {\it supertrace} if
$$str(f \cdot g)=(-1)^{\pi(f)\pi(g)}str(g \cdot f) \ \mbox{ for all } f,g\in {\mathcal A}.$$
\end{definition}

\begin{definition}\label{tr}
A linear function $tr$ on ${\mathcal A}$ is called a {\it trace} if
$$tr(f \cdot g)=tr(g \cdot f) \ \mbox{ for all } f,g\in {\mathcal A}.$$
\end{definition}

We will use the notation $\str$ and the term "(super)trace" to denote both cases,
traces and supertraces, simultaneously.

\section{The superalgebra of observables}

Let $V={\mathbb  R}^N$ be endowed with a positive definite symmetric
bilinear form $(\cdot,\cdot)$.
For any nonzero $\vv \in V$, define the {\it
reflections} $r_\vv$ as follows:
\be\label{ref}
r_\vv: \ \ \
{\vec x} \mapsto {\vec x} -2 \frac {({\vec x},\,\vv)} {(\vv,\,\vv)} \vv \qquad
\mbox{ for any }{\vec x} \in V.
\ee

A finite set of non-zero vectors $\R\subset V$ is said to be a {\it root system}
and any vector $\vv\in \R$ is called a \emph{root}
if the following conditions hold:

i) $\R$ is ${r}_{\vec w}$-invariant for any $\vec w \in \R$,

ii) if $\vv_1,\vv_2\in \R$ are proportional to each other, then either $\vv_1=\vv_2$ or $\vv_1=-\vv_2$.

The Coxeter group $G\subset O(N, {\mathbb  R})\subset End(V)$ generated
by all reflections ${r}_\vv$ with $\vv \in \R$ is finite.

We do not apply any conditions on the scalar products of the roots because we
want to consider
both crystallographic and non-crystallographic root systems, e.g., $I_2(n)$ (see Theorem \ref{th2}).

Let $\eta$ be a complex-valued $G$-invariant function on $\R$, i.e.,
$\eta(\vv)=\eta({\vec w})$ if $r_\vv$ and $r_{\vec w}$ belong to
one conjugacy class of $G$.

We consider here
the Symplectic Reflection (Super)algebra
over complex numbers
(see \cite{sra}) $H:=H_{1,\eta}(G)$
and call it the
\emph{superalgebra of observables of Calogero model based on root system $\R$}.
\footnote
{This algebra
has a faithful
representation via Dunkl differential-difference operators $D_i$,
see \cite{Dunkl}, acting on the space of $G$-invariant smooth functions on $V$,
namely $\hat a^{\alpha}_i =\frac 1
{\sqrt{2}} (x_i + (-1)^\alpha D_i)$, see \cite{BHV,Poly}.
The Hamiltonian of the Calogero model based on the root system
\cite{Cal1, Cal2, Cal3, OP} is the operator $\hat T^{01}$  defined in (\ref{sl2})
(see \cite{BHV} ).
The wave functions are obtained in this model
via the standard Fock procedure with the Fock vacuum $ |0\rangle$ such that
$\hat a_i^0 |0\rangle$=0 for all $i$ by acting on $ |0\rangle$ with $G$-invariant polynomials
of the~$\hat a^1_i$.
}

This algebra consists of noncommuting polynomials in $2N$ indeterminates
$a^\alpha_i$, where $\alpha=0,1$ and $i=1,\, ... ,\, N$, with coefficients
in $\mathbb C[G]$ satisfying the relations (see \cite{sra} Eq. (1.15))%
\footnote{
The sign and coefficient of the sum in the rhs of Eq. (\ref{comaa})
is chosen for obtaining the Calogero model in the form
\cite{BHV}, Eq. (1), Eq. (5),  Eq. (9), Eq. (10) when $\R$ is of type
$A_{N-1}$.
}
\be\label{comaa}
[a^{\alpha}_i, a^{\beta}_j] = \varepsilon^{\alpha\beta}
\left(\delta_{ij}+
\sum_{\vv\in\R} \eta(\vv) \frac {v_i v_j}{(\vv,\,\vv)}r_{\vv}\right),
\ee
and
\be\label{comav}
r_{\vv} a^{\alpha}_i = \sum_{j=1}^N \left(\delta_{ij} - 2
\frac {v_i v_j}{(\vv,\,\vv)}\right)a^{\alpha}_j  r_{\vv}.
\ee

Here $\varepsilon^{\alpha\beta}$ is the antisymmetric tensor such that $\varepsilon^{01}=1$,
and $v_i$ ($i=1,..., N$) are the coordinates of the vector $\vv$.
The commutation relations (\ref{comaa}), (\ref{comav}) suggest
to define the {\it parity} $\pi$ by setting:
\be
\pi (a^{\alpha}_i)=1
\ \mbox{ for any }\alpha,\ i;
\qquad \pi( r_{\vv})=0 \ \mbox{ for any } \vv \in \R.
\ee
and we can consider the algebra $H$ as a superalgebra as well.


\section{$\mathfrak {sl}_2$}

Observe an important property of the superalgebra
$H$: The Lie (super)algebra of its inner
derivations
contains the Lie subalgebra $\mathfrak{sl}_2$ generated by operators
\be
D^{\alpha\beta}: \ \ \ f\mapsto D^{\alpha\beta}f=[T^{\alpha\beta}, \, f]
\ee
where $\alpha,\beta=0,1$, and $f\in H$, and polynomials $T^{\alpha\beta}$ are defined
as follows:
\be \label {sl2}
T^{\alpha\beta}:= \frac 1 2 \sum _{i=1}^N
\left( a^{\alpha}_i a^{\beta}_i + a^{\beta}_i a^{\alpha}_i \right )\, .
\ee
These operators satisfy the following relations:
\be \label{csl2}
[D^{\alpha\beta}, D^{\gamma\delta}]= \epsilon^{\alpha\gamma}
D^{\beta\delta} +\epsilon^{\alpha\delta} D^{\beta\gamma} +
\epsilon^{\beta\gamma} D^{\alpha\delta} +\epsilon^{\beta\delta}
D^{\alpha\gamma}\,,
\ee
since
\[
[T^{\alpha\beta}, T^{\gamma\delta}]= \epsilon^{\alpha\gamma}
T^{\beta\delta} +\epsilon^{\alpha\delta} T^{\beta\gamma} +
\epsilon^{\beta\gamma} T^{\alpha\delta} +\epsilon^{\beta\delta}
T^{\alpha\gamma}.
\]
It follows from Eq. (\ref{csl2}) that
the operators $D^{00}$, $D^{11}$ and $D^{01}=D^{10}$
constitute an $\mathfrak{sl}_2$-triple:
$$[D^{01},\, D^{11}] = 2 D^{11},\qquad
[D^{01},\, D^{00}] = - 2 D^{00},\qquad
[D^{11},\, D^{00}] =  - 4 D^{01}.$$

The polynomials $T^{\alpha\beta}$ commute with ${\mathbb C}[G]$, i.e.,
$[T^{\alpha\beta},\,r_\vv]=0$,
and act on the $a^{\alpha}_i$ as on vectors
of the irreducible 2-dimensional
$\mathfrak{sl}_2$-modules:
\be\label{sl2vec}
D^{\alpha\beta}a^{\gamma}_i=\left[T^{\alpha\beta},\,a^{\gamma}_i\right]=
\varepsilon^{\alpha\gamma}a^{\beta}_i +
\varepsilon^{\beta\gamma}a^{\alpha}_i, \quad \mbox{ where } i=1,\,\dots\,, N.
\ee

We will denote this $\mathfrak{sl}_2$ thus realized by the symbol $SL2$.

The subalgebra
\be
H_0:=\{f\in H \mid D^{\alpha\beta}f=0 \mbox { for any } \alpha,\, \beta \}\subset H
\ee
is called the \emph{subalgebra of singlets}.

Introduce also the subspaces $H_s:=\oplus_{i_s=1}^\infty H^{i_s}_s$,
which is the direct sum of all irreducible $SL2$-modules $H^{i_s}_s$  of
 spin $s$, for $s=0,\, 1/2,\, 1,\,...$. It is clear that $H_0$ is
 the defined above subalgebra of singlets.

The (super)algebra $H$ can be decomposed in the following way
\[
H=H_0 \oplus H_{rest},  \text{\ \ where \ \ }
H_{rest}:= H_{1/2} \oplus H_1 \oplus H_{3/2} \oplus\dots.
\]
Then each element $f\in H$
can be represented in the form $f=f_0+f_{rest}$, where $f_0 \in H_0$ and $f_{rest}\in H_{rest}$.

Note, that since $SL2$ is generated by inner derivations and $T^{\alpha\beta}$ are even elements,
each two-sided ideal $\mathcal I\subset H$ can be decomposed in an analogous way:
$\mathcal I=\mathcal I_0\oplus \mathcal I_{1/2}\oplus... $.

Since $T^{\alpha\beta}$ are even elements of the superalgebra $H$,
we have
$\str(D^{\alpha\beta}f)=0$ for any (super)trace $\str$ on $H$, and hence
the following proposition takes place%
\footnote{This elementary fact is known for a long time, see, eg, \cite{KV}.}:
\begin{proposition}\label{f0}
$\str(f)=\str(f_0)$ for any $f\in H$ and any (super)trace $\str$ on $H$.
\end{proposition}

\begin{proof}
If $s\ne 0$, then the elements of the form $D^{\alpha\beta} f $ , where
$\alpha,\, \beta =0,\,1$, and $f \in H^{i_s}_s$, $f\ne0$, span the irreducible $SL2$-module $H^{i_s}_s$.
This implies $\str f =0$ for any (super)trace on $H$ and any $f\in H_{rest}$.
\end{proof}


\section{The (super)traces on $H$}\label{trace}

It is shown in \cite{KV,KTSRA,KTroot} that the algebra $H$ has a multitude
of independent (super)traces. For the list of dimensions of the spaces of the
(super)traces
on $H_{1,\eta}(M)$ for all finite Coxeter groups $M$, see \cite{stek}.
In particular,  there is an $m$-dimensional space of
traces and an $(m+1)$-dimensional space of supertraces on $H_{1,\eta}(I_2(2m+1))$.

Every (super)trace $\str (\cdot)$ on any associative (super)algebra ${\mathcal A}$ generates
the following bilinear form on ${\mathcal A}$:
\bee\label{bf2}
B_{\str }(f,g)=\str (f\cdot g) \mbox{ for any } f,g\in {\mathcal A}.
\eee

It is obvious that if such a bilinear form $B_{\str }$ is degenerate,
then
the
kernel of this form (i.e., the set of all vectors $f \in {\mathcal A}$ such that $B_{\str}(f, g)=0$ for any
$g\in {\mathcal A}$)
is the two-sided
ideal ${\mathcal I^{\str}}\subset {\mathcal A}$.

The ideals of this sort are found, for example,
in \cite[Theorem 9.1]{2m+1}  (generalizing the results of \cite{V1, V2}  and \cite{K2} for the two-
and three-particle Calogero models).

Theorem 9.1 from \cite{2m+1} may be shortened to the following theorem:
\vskip 2mm
\begin{theorem}\label{th2}
Let $ m\in \mathbb Z$, where $m\geqslant 1$
and $n=2m+1$.
Then

\emph{1)} The associative algebra $H_{1,\eta}(I_2(n))$
has  nonzero  traces $tr_\eta$ such that the symmetric invariant bilinear form
$B_{tr_\eta}(x,y)=tr_\eta(x\cdot y)$  is degenerate if and only if
$\eta= \frac z n$, where $z\in \mathbb Z \setminus n\mathbb Z$.
All such traces are proportional to each other.

\emph{2)} The associative superalgebra $H_{1,\eta}(I_2(n))$
has  nonzero supertraces $str_\eta$ such that the supersymmetric invariant bilinear form
$B_{str_\eta}(x,y)=str_\eta(x\cdot y)$  is degenerate if
$\eta= \frac z n$, where $z\in \mathbb Z \setminus n\mathbb Z$.
All such supertraces are proportional to each other.

\emph{3)} The associative superalgebra $H_{1,\eta}(I_2(n))$
has  nonzero supertraces $str_{\eta}$ such that the supersymmetric invariant bilinear form
$B_{str_{\eta}}(x,y)=str_{\eta}(x\cdot y)$  is degenerate if
$\eta= z + \frac 1 2$, where $z\in \mathbb Z$.
All such supetraces are proportional to each other.

\emph{4)} For all other values of $\eta$, all nonzero traces and supertraces are nondegenerate.
\end{theorem}

Theorem \ref{th2} implies that if $z\in \mathbb Z \setminus n\mathbb Z$, then
there exists the degenerate trace $tr_z$ generating the ideal $\mathcal I^{tr_z}$ consisting of the kernel
of the degenerate form $B_{tr_z}(f,g)=tr_z(f\cdot g)$, and simultaneously
the degenerate supertrace $str_z$ generating the ideal $\mathcal I^{str_z}$ consisting of the kernel
of the degenerate form $B_{str_z}(f,g)=str_z(f\cdot g)$.

A question arises: is it true that $\mathcal I^{tr_z}=\mathcal I^{str_z}$?

Answer to this and other similar questions
can be considerably simplified by considering only the singlet parts of these ideals.

The following theorem justifies this method:

\vskip 2mm
\begin{theorem}\label{th3}
Let $\str_1$ and $\str_2$ be degenerate (super)traces on
$H$.
They generate the
two-sided ideals $\mathcal I_1$ and $\mathcal I_2$ consisting of the kernels
 of bilinear forms $B_1 (f,g)=\str_1(f\cdot g)$ and
$B_2 (f,g)=\str_2(f\cdot g)$, respectively.

Then
$\mathcal I_{1}=\mathcal I_{2}$ if and only if
$\mathcal I_{1} \bigcap H_0 = \mathcal I_{2} \bigcap H_0$.
\end{theorem}

\begin{proof}
It suffices to prove that
if
$\mathcal I_{1} \bigcap H_0 = \mathcal I_{2} \bigcap H_0$,
then
$\mathcal I_{1}=\mathcal I_{2}$.

Consider any non-zero element $f\in \mathcal I_1$.
For any $g\in H$, we have $\str_1(f\cdot g)=0$,
$f\cdot g\in \mathcal I_1$ and $(f\cdot g)_0\in \mathcal I_1$.
So
$(f\cdot g)_0\in \mathcal I_1 \bigcap H_0$.
Due to hypotheses of this Theorem,
$(f\cdot g)_0\in \mathcal I_2 \bigcap H_0$,
and hence ${\str_2((f\cdot g)_0)=0}$.
 Proposition \ref{f0} gives
$\str_2(f\cdot g)=\str_2((f\cdot g)_0)$
which implies
$\str_2(f\cdot g)=0$.

Therefore, $f\in \mathcal I_2$.
\end{proof}


\section*{Acknowledgments}
The authors (S.K. and I.T.) are grateful to Russian Fund for Basic Research
(grant No.~${\text{17-02-00317}}$)
for partial support of this work.


\end{document}